OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = 2x2 – 8x + 2 a. (-4, 66) b. (4, 2) c. (-2, 26) d. (2, -6)

OpenStudy (anonymous):

(2,-6)

OpenStudy (anonymous):

when you get the equation in std form, the x-coordinate of the vertex can be found by the following formula \[x=\frac{-b}{2a}\]where your quadratic is\[y=ax^2+bx+c\]

OpenStudy (anonymous):

Oh ,, okay thanks , so thats the formual i have ta use when finding the vertex ?

OpenStudy (anonymous):

It just gives you the x-coordinate of the vertex. To get the y-coordinate, plug the value of x you got back into the equation of the parabola. You only use this formula when the equation is given in the y=ax^2+bx+c form.

OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = 7x2 + 21x – 1 Question 3 answers a. (-2, -15) b. (2, 69) c. (1, 27) d. (-3/2, -67/4)

OpenStudy (anonymous):

ok, your turn

OpenStudy (anonymous):

Vertex of y=ax^2+bx+c is \( \large \left(\frac{-b}{2a}, \frac{-b^2+4ac}{4a} \right)\)